A ticket on the roller coaster is priced differently for adults and children. One day there were 5 adults and 8 children in a group and the cost of their tickets was 48.50. Another group of 4 adults and 12 children paid 57. What is the price of each type of ticket?
I tried to do 9a+20c=105.5
a+c=29

Answer by Greenfinch(383)

Let adult cost be A and childrens' cost be C
Then 5A + 8C = 4850
and 4A + 12C = 5700
Multiply 1 by 3 and 2 by 2
15A + 24C = 14550
8A + 24C = 11400 so, subtracting
7A = 3150
A = 450 or $4.50
Substituting in 1
5 x 450 + 8C = 4850
2250 +8C = 4850
8C = 2600
C = 325
So Adult tickets cost $4.50 and Childrens' tickets cost £3.25